# Environmentally Sustainable Software Design and Development: A Systematic Literature Review

### v1

@@ Acknowledgements.
- Wedyan et al. (2023) Fadi Wedyan, Rachael Morrison, and Osama Sam Abuomar. 2023. Integration and Unit Testing of Software Energy Consumption. In Tenth International Conference on Software Defined Systems, SDS 2023, San Antonio, TX, USA, October 23-25, 2023, Muhannad Quwaider, Yaser Jararweh, Elhadj Benkhelifa, and Jaime Lloret (Eds.). IEEE, 60–64. https://doi.org/10.1109/SDS59856.2023.10329262
- Wei et al. (2023) Xiaokai Wei, Sujan Kumar Gonugondla, Shiqi Wang, Wasi Uddin Ahmad, Baishakhi Ray, Haifeng Qian, Xiaopeng Li, Varun Kumar, Zijian Wang, Yuchen Tian, Qing Sun, Ben Athiwaratkun, Mingyue Shang, Murali Krishna Ramanathan, Parminder Bhatia, and Bing Xiang. 2023. Towards Greener Yet Powerful Code Generation via Quantization: An Empirical Study. In Proceedings of the 31st ACM Joint European Software Engineering Conference and Symposium on the Foundations of Software Engineering, ESEC/FSE 2023, San Francisco, CA, USA, December 3-9, 2023, Satish Chandra, Kelly Blincoe, and Paolo Tonella (Eds.). ACM, 224–236. https://doi.org/10.1145/3611643.3616302
- Wohlin (2014) Claes Wohlin. 2014. Guidelines for snowballing in systematic literature studies and a replication in software engineering. In 18th International Conference on Evaluation and Assessment in Software Engineering, EASE ’14, London, England, United Kingdom, May 13-14, 2014, Martin J. Shepperd, Tracy Hall, and Ingunn Myrtveit (Eds.). ACM, 38:1–38:10. https://doi.org/10.1145/2601248.2601268
@@
- Zulfiqar et al. (2023) Muhammad Zulfiqar et al. 2023. Digitalized world and carbon footprints: does digitalization really matter for sustainable environment? Environ. Sci. Pollut. Res. 30 (2023), 88789–88802. https://doi.org/10.1007/s11356-023-28332-z

### v2

@@ Acknowledgements.
- Wedyan et al. (2023) Fadi Wedyan, Rachael Morrison, and Osama Sam Abuomar. 2023. Integration and Unit Testing of Software Energy Consumption. In Tenth Int. Conf. on Software Defined Systems, SDS 2023, San Antonio, TX, USA, October 23-25, 2023, Muhannad Quwaider et al. (Eds.). IEEE, 60–64. https://doi.org/10.1109/SDS59856.2023.10329262
- Wei et al. (2023) Xiaokai Wei, Sujan Kumar Gonugondla, Shiqi Wang, Wasi Uddin Ahmad, Baishakhi Ray, Haifeng Qian, Xiaopeng Li, Varun Kumar, Zijian Wang, Yuchen Tian, Qing Sun, Ben Athiwaratkun, Mingyue Shang, Murali Krishna Ramanathan, Parminder Bhatia, and Bing Xiang. 2023. Towards Greener Yet Powerful Code Generation via Quantization: An Empirical Study. In Proceedings of the 31st ACM Joint European Software Engineering Conf. and Symposium on the Foundations of Software Engineering, ESEC/FSE 2023, San Francisco, CA, USA, December 3-9, 2023, Satish Chandra et al. (Eds.). ACM, 224–236. https://doi.org/10.1145/3611643.3616302
- Wohlin (2014) Claes Wohlin. 2014. Guidelines for snowballing in systematic literature studies and a replication in software engineering. In 18th International Conference on Evaluation and Assessment in Software Engineering, EASE ’14, London, England, United Kingdom, May 13-14, 2014, Martin J. Shepperd et al. (Eds.). ACM, 38:1–38:10. https://doi.org/10.1145/2601248.2601268
@@
- Zulfiqar et al. (2023) Muhammad Zulfiqar, Safdar Husain Tahir, Muhammad Rizwan Ullah, and Sadeen Ghafoor. 2023. Digitalized world and carbon footprints: does digitalization really matter for sustainable environment? Environ. Sci. Pollut. Res. 30 (2023), 88789–88802. https://doi.org/10.1007/s11356-023-28332-z